OpenStudy (anonymous):

Michael takes 90 min to type 5850 words. If he increases his typing speed by 5 words/min, how many minutes will he take to type 10,500 words?

OpenStudy (anonymous):

Help?

OpenStudy (anonymous):

How many words/minute can he type before increasing it?

OpenStudy (anonymous):

5?

OpenStudy (anonymous):

No..take the words divided by number of minutes 5850/90=?

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

what is it?

OpenStudy (anonymous):

65

OpenStudy (anonymous):

Yeah now thats the words per minute. increase it by 5

OpenStudy (anonymous):

multiply

OpenStudy (anonymous):

325

OpenStudy (anonymous):

What is 65+5?

OpenStudy (anonymous):

13

OpenStudy (anonymous):

I think you should brush up on basic math before you do a problem like this.

OpenStudy (anonymous):

excuse me ?

OpenStudy (anonymous):

but that is not a choice

OpenStudy (anonymous):

cant i make a proportion out of this

OpenStudy (anonymous):

90 minutes to type 5850 words. If she increased her typing speed by 5 words/minute.....you have 90 minutes, increase by 5 words per minute...90 x 5 = 450. So she has increased her words by 450 which equals (5850 + 450) = 6300. proportion time... 90/6300 = x/10,500 (cross multiply) (6300)(x) = (90)(10,500) 6300x = 945000 (divide by 6300 on both sides) 6300x/6300 = 945000/6300 x = 150 ANSWER : It would take her 150 minutes to type 10,500 words

OpenStudy (anonymous):

ok thank u u really r a life saver

OpenStudy (anonymous):

I hope I did that right....I am looking at CalebBeavers solution...let me see if his answer will be the same when worked out.

OpenStudy (anonymous):

g2g to sleep

OpenStudy (anonymous):

bye

OpenStudy (anonymous):

ut do it n ill check it tomorrow

OpenStudy (anonymous):

bye

OpenStudy (anonymous):

90/5850 = 1/x 90x = 5850 x = 65 increased by 5 words per minute = 70 1/70 = x/10,500 70x = 10500 x = 10500/70 x = 150 Thank God.....the same answer either way.

OpenStudy (anonymous):

ok